OpenStudy (anonymous):

Solve the equation: 3^(1-4y)=sqrt3

myininaya (myininaya):

1-4y=1/2 -4y=1/2-1 -4y=-1/2 4y=1/2 y=4/2 y=2

myininaya (myininaya):

oops y=1/2*1/4=1/8

myininaya (myininaya):

you see 4y=1/2 i accidently divided lhs by 4 while i multiplied the other side by 4 thats a big mistake

myininaya (myininaya):

do you see this?

OpenStudy (anonymous):

OK, just equating exponents.

OpenStudy (anonymous):

I see what you did myininaya. Thank you!

myininaya (myininaya):

cool!